### v1

# ControlFusion: A Controllable Image Fusion Framework with Language-Vision Degradation Prompts

Abstract
Current image fusion methods struggle to address the composite degradations encountered in real-world imaging scenarios and lack the flexibility to accommodate user-specific requirements. In response to these challenges, we propose a controllable image fusion framework with language-vision prompts, termed ControlFusion, which adaptively neutralizes composite degradations. On the one hand, we develop a degraded imaging model that integrates physical imaging mechanisms, including the Retinex theory and atmospheric scattering principle, to simulate composite degradations, thereby providing potential for addressing real-world complex degradations from the data level. On the other hand, we devise a prompt-modulated restoration and fusion network that dynamically enhances features with degradation prompts, enabling our method to accommodate composite degradation of varying levels. Specifically, considering individual variations in quality perception of users, we incorporate a text encoder to embed user-specified degradation types and severity levels as degradation prompts. We also design a spatial-frequency collaborative visual adapter that autonomously perceives degradations in source images, thus eliminating the complete dependence on user instructions. Extensive experiments demonstrate that ControlFusion outperforms SOTA fusion methods in fusion quality and degradation handling, particularly in countering real-world and compound degradations with various levels.
Figure 1: Fusion performance comparison across real-world, composite degradation, and varying degradation levels.
1 Introduction
Image fusion is a crucial technique in image processing. By effectively leveraging complementary information from multiple sources, the limitations associated with information collection by single-modal sensors can be significantly mitigated [40]. Among the various research areas within this field, infrared-visible image fusion (IVIF) has garnered considerable attention. IVIF integrates essential thermal information from infrared (IR) images with the intricate texture details from visible (VI) images, offering a richer and more complete depiction of the scene [44]. By harmonizing diverse information and producing visually striking results, IVIF has found extensive applications in areas such as security surveillance [45], military detection [23], scene understanding [7, 43], and assisted driving [1]etc.
Recently, IVIF has emerged as a focal point of research, leading to rapid advancements in related methods. Based on the architecture of deep networks, these methods can be categorized into convolutional neural network- [14, 50], autoencoder- [11, 12], generative adversarial network- [19, 15], Transformer- [21, 42], and diffusion model-based [49, 41] methods. Besides, from a functional perspective, they can be categorized into visual-oriented [30, 50], joint registration-fusion [28, 37], semantic-driven [29, 16], and degradation-robust [32, 41] schemes.
Notably, although degradation-robust schemes can mitigate degradation interference to some extent, they still have certain limitations. On the one hand, these methods often rely on simplistic approaches to construct training data, overlooking the domain gap between simulated data and realistic images, which significantly hampers their generalizability in practical scenarios, as presented in Fig. 1(I). On the other hand, they are tailored for specific or single types of degradation, making them ineffective in handling more complex composite degradations, as illustrated in Fig. 1(II). Last but not least, as shown in Fig. 1(III), existing methods lack degradation level modeling, causing a sharp decline in performance as degradation intensifies. Furthermore, these approaches do not offer the flexibility to dynamically adjust fusion results according to the diverse perceptual requirements of different users.
To overcome these limitations, we propose a versatile and controllable fusion model based on language-vision prompts, termed ControlFusion. On the one hand, we introduce a physics-driven degraded imaging model, which accurately simulates the physical mechanisms of the actual degradation process, e.g., illumination-, weather-, and sensor-related distortions. The proposed model not only effectively bridges the gap between synthetic data and real-world images but also offers valuable data support for handling complex composite degradations. On the other hand, we develop a prompt-modulated image restoration and fusion network to generate high-quality fusion results. The prompt-modulated module enables our network to dynamically adjust feature distribution based on degradation characteristics (which can be specified by users), achieving robust feature enhancement. This also allows our method to respond to diverse user requirements. Furthermore, we devise a spatial-frequency visual adapter that combines frequency-domain degradation priors to directly extract degradation cues from degraded inputs, eliminating the heavy reliance on user instructions, as customizing prompts for each scene is time-consuming and labor-intensive. As shown in Fig. 1, benefiting from the aforementioned designs, our method excels in handling both real-world and composite degradation scenarios, and it shows strong robustness across various degradation levels. In summary, our main contributions are as follows:
    We propose ControlFusion, a versatile image restoration and fusion framework that uniformly models diverse degradation types and degrees, using textual and visual prompts as a medium. Specifically, its controllability enables it to respond to user-specific customization needs.
    A spatial-frequency visual adapter is devised to integrate frequency characteristics and directly extract text-aligned degradation prompts from visual images, enabling large-scale automated deployment.
    A physics-driven imaging model is developed, integrating physical mechanisms such as the Retinex theory and atmospheric scattering principle to bridge the gap between synthetic data and real-world images, while providing data support for addressing composite degradations.
2 Related Work
Image Fusion. Earlier visual-oriented fusion approaches primarily concentrated on merging complementary information from multiple modalities and improving visual fidelity. These methods often depend on sophisticated network architectures and carefully designed loss functions to ensure that the integrated content retains the essential characteristics of the original images, preserving their unique features while enhancing overall quality. The mainstream network architectures primarily include CNN-based [20, 47], AE-based [12, 48], GAN-based [19, 15], Transformers [21] and diffusion models [49]. Furthermore, several schemes including joint registration and fusion [33, 37], semantic-driven [29, 15], and degradation-robust [32, 38, 41] methods, are proposed to broaden the practical applications of image fusion. For instance, Tang et al. [31] and Liu et al. [17] developed corresponding solutions for illumination distortions and noise interference, respectively. In addition, Yi et al. [38] proposed a unified image restoration and fusion network to address various degradations. However, they are unable to handle mixed degradations and struggle to generalize to real-world scenarios. In particular, Text-IF [38] relies on tedious manual efforts to customize text prompts for each scene, hindering large-scale automated deployment.
Image Restoration with Vision-Language Models. With the advancement of deep learning toward multimodal integration, the image restoration field has progressively evolved into a new paradigm of text-driven schemes. CLIP [26] establishes visual-textual semantic representation alignment through dual-stream Transformer encoders pre-trained on 400 million image-text pairs. This framework lays the foundation for the widespread adoption of Prompt Engineering in computer vision. Recent studies integrate CLIP encoders with textual user instructions, proposing generalized restoration frameworks including PromptIR [24], AutoDIR [9], and InstructIR [2], which effectively handle diverse degradation types. Moreover, SPIRE [25] further introduces fine-grained textual restoration cues by quantifying degradation severity levels and supplementing semantic information for precision restoration. To eliminate reliance on manual guidance, Luo et al. [18] developed DA-CLIP by fine-tuning CLIP on mixed-degradation datasets, enabling degradation-aware embeddings to facilitate distortion handling. However, existing unified restoration methods are primarily designed for natural images, which exhibit notable limitations when processing multimodal image pairs (e.g., infrared-visible).
Figure 2: The overall framework of our semantic and degradation dual-prior guided image fusion network.
3 Physics-driven Degraded Imaging Model
3.1 Imaging Formulation
In real-world scenarios, images are often affected by multiple degradation factors simultaneously, which leads to suboptimal performance of restoration and fusion methods designed for isolated degradations. To tackle the complex and variable degradation challenges, we propose a physics-driven imaging model aimed at reducing the domain gap between simulated data and real-world imagery. Visible (VI) images are typically degraded by illumination conditions (low-light, overexposure), weather (rain, haze), and sensor-related issues (noise, blur), while infrared (IR) images usually suffer from sensor-related interference, such as stripe noise and low contrast. Thereby, given a clear image $I_{m}(m\in\{ir,vi\})$, the proposed imaging model can be mathematically represented as:
$D_{m}=\mathcal{P}_{s}\left(\mathcal{P}_{w}\left(\mathcal{P}_{i}\left(I_{m}\right)\right)\right),$ (1)
where $D_{m}$ is the corresponding degraded image, and $\mathcal{P}_{i}$, $\mathcal{P}_{w}$, and $\mathcal{P}_{s}$ represent illumination-, weather-, andsensor-related distortions, respectively.
Illumination-related Distortions. Following the theoretical framework of Retinex, the formation of illumination-degraded images $D^{i}_{vi}$ is mathematically modeled as:
${D^{i}_{vi}=\mathcal{P}_{i}(I_{vi})=\frac{I_{vi}}{L}\cdot L^{\gamma},}$ (2)
where $L$ is the illumination map estimated by LIME [5]. $\gamma\in[0.5,3]$ controls the illumination level.
Weather-related Distortions. According to the methodologies in [22, 13], we employ the following formula to simulate weather-related degradations (i.e. rain and haze):
$D^{w}_{vi}=\mathcal{P}_{w}(I_{vi})=I_{vi}\cdot t+A(1-t)+R,$ (3)
where $t$ and $A$ denote the transmission map and atmospheric light, respectively. And $t$ is defined by the exponential decay of light, expressed as $t=e^{-\beta d(x)}$, where the haze density coefficient $\beta\in[0.5,2.0]$. $d(x)$ refers to the scene depth, estimated by DepthAnything-V2 [39]. Besides, $A$ is constrained within $[0.3,0.9]$ for realistic simulation.
Sensor-related Distortions. Sensor-related distortions encompass various types of noise, motion blur, and contrast degradation. Contrast degradation and stripe noise of infrared images can be modeled as:
$D^{s}_{ir}=\mathcal{P}_{s}(I_{ir})=\alpha\cdot I_{ir}+\mathbf{1}_{H}\mathbf{n}^{\top},$ (4)
where $\alpha$ is a constant less than 1 for contrast reduction, $\mathbf{1}_{H}\in\mathbb{R}^{H}$ is an all-ones column vector and $\mathbf{n}\in\mathbb{R}^{W}$ represents the column-wise Gaussian noise vector sampled from $\mathcal{N}(0,\epsilon^{2})$ with $\epsilon\in[1,15]$. Moreover, Gaussian noise and motion blur in source images can be modeled as:
$D^{s}_{m}=\mathcal{P}_{s}(I_{m})=I_{m}\ast K(N,\theta)+\mathcal{N}(0,\sigma^{2}),$ (5)
where $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise with $\sigma\in[5,20]$. * denotes convolution with a blur kernel $K(N,\theta)=\frac{1}{N}R_{\theta}(\delta_{c}\otimes\mathbf{1}_{N})$ constructed by rotating an impulse $\delta_{c}\otimes\mathbf{1}_{N}$ with random angle $R_{\theta}$ ($\theta\in[10,80]$), using $\frac{1}{N}$ to normalize the kernel energy. Here, $N\in[3,12]$ controls the blur level.
3.2 Composite Degradation Dataset with Levels
Based on Eq. (1), we construct a multi-modal composite Degradation Dataset with four Levels (DDL-12), comprising 12 distinct degradations along with their corresponding high-quality images. In particular, we selecte $2,050$ high-quality clear images from multiple publicly available infrared-visible datasets [36, 16, 8, 30]. Subsequently, the degraded imaging model is employed to synthesize degraded images with 12 types and 4 levels of degradation. Finally, the DDL-12 dataset contains approximately $32,000$ training image pairs and $4,000$ test image pairs.
4 Methodology
4.1 Problem Formulation
Given two source images $I_{ir}\in\mathbb{R}^{H\times W\times 1}$ and $I_{vi}\in\mathbb{R}^{H\times W\times 3}$, the typical fusion paradigm employs a deep network $\mathcal{N}_{f}$ to integrate complementary information and synthesize the fused image $I_{f}$, expressed as:
$I_{f}=\mathcal{N}_{f}(I_{ir},I_{vi}).$ (6)
However, in complex scenarios, source images usually suffer from degradation interference, thus the advanced fusion paradigm must take both image restoration and information aggregation into account. Concatenation and coupling are two common approaches to implement this. Notably, while the concatenation approach is straightforward, it neglects the intrinsic relationship between restoration and fusion, frequently yielding suboptimal outcomes. As illustrated in Fig. 2, we propose a controllable paradigm that couples restoration and fusion with degradation prompt, termed ControlFusion, to overcome this limitation. On the one hand, the coupled approach enhances task synergy. On the other hand, leveraging degradation prompts to modulate the restoration and fusion process enables the network to adapt to diverse degradation distributions while meeting user-specific customization requirements. Thus, the proposed restoration and fusion paradigm can be defined as:
$I_{f}=\mathcal{N}_{rf}(I_{ir},I_{vi},p~{}|~{}\Omega),$ (7)
where $\mathcal{N}_{rf}$ is a restoration and fusion network, $P$ denotes degradation prompts, and $\Omega$ indicates the degradation set. Notably, degradation prompts can be either user-specified text or directly extracted from input images, greatly enhancing the flexibility of our framework without heavily relying on user instructions. To fully harness its potential, our ControlFusion employs a two-stage optimization protocol. Stage I aligns textual prompts with visual embedding, while Stage II optimizes the overall framework.
4.2 Stage I: Textual-Visual Prompts Alignment
Text Semantic Encoder. Adopting text to represent the degradation type and degree is the most straightforward approach. Thus, we introduce a pre-trained vision-language model (i.e., CLIP [26]), to achieve prompt embedding. Given a user instruction $T_{instr}$, the text semantic encoder $\mathcal{E}_{t}$ from CLIP is deployed to convert it into a text embedding:
$p_{text}=\mathcal{E}_{t}(T_{instr}).$ (8)
Figure 3: The visualization of various types of degradation in the spatial and frequency domains.
Spatial-Frequency Collaborative Visual Adapter. It is important to emphasize that customizing text instructions for each image is extremely time-consuming and labor-intensive. Therefore, we expect to develop a textual-visual adapter that directly extracts visual embeddings from source images that align with $p_{text}$, enabling automated deployment. In particular, as shown in Fig. 2, images with different degradations in the same scene exhibit distinct spectral characteristics, indicating that frequency features contain rich degradation priors. To this end, we devise a spatial-frequency collaborative visual adapter (SFVA) to extract visual prompts. In the frequency branch, we utilize the Fast Fourier Transform (FFT) algorithm in conjunction with a CNN to extract frequency features $F_{fre}$:
$F_{fre}^{m}$ $=\textstyle{\sum_{x=0}^{W-1}\sum_{y=0}^{H-1}D_{m}(x,y)e^{-j2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)},}$ (9)
$F_{fre}$ $=\text{Linear}\left(\text{Conv}\left([F_{fre}^{ir},F_{fre}^{vi}]\right)\right),$
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes channel-wise concatenation. In the spatial branch, we employ cropping and downsampling for data augmentation, followed by a CNN-based visual encoder to extract spatial features $F_{spa}$. Then, we concatenate $F_{fre}$ and $F_{spa}$, and apply successive linear layers to obtain the final visual embedding $p_{vis}$. Furthermore, we employ MSE loss $\mathcal{L}_{mse}$ and cosine similarity loss $\mathcal{L}_{cos}$ to impose constraints on the similarity between $p_{vis}$ and the pre-defined $p_{text}$. The loss for Stage I is formulated as:
$\mathcal{L}_{I}=\lambda_{1}\underbrace{\|p_{vis}-p_{text}\|^{2}}_{\mathcal{L}_{\text{mse}}}+\lambda_{2}\underbrace{(1-\frac{p_{vis}\cdot p_{text}}{\|p_{vis}\|\|p_{text}\|})}_{\mathcal{L}_{\text{cos}}},$ (10)
where $\lambda_{1}$ and $\lambda_{2}$ are parameters, balancing various losses.
4.3 Stage II: Prompt-modulated Restoration and Fusion
4.3.1 Network Architectures
Feature Encoding and Fusion Layer. As shown  Fig. 2, we devise the hierarchical transformer-based encoders to extract multi-scale feature representations from degraded infrared ($D_{ir}$) and visible ($D_{vi}$) images separately, which is formulated as:
$\{F_{ir},F_{vi}\}=\{\mathcal{E}_{ir}(D_{ir}),\mathcal{E}_{vi}(D_{vi})\},$ (11)
where $\mathcal{E}_{ir}$ and $\mathcal{E}_{vi}$ are infrared and visible image encoders.
Furthermore, to achieve comprehensive cross-modal feature integration, we design the intra-domain fusion unit, where the cross-attention (CR-ATT) mechanism is employed to facilitate interaction between features across modalities. Specifically, the linear transformation functions $\mathcal{F}_{ir}^{qkv}$ and $\mathcal{F}_{vi}^{qkv}$ project $F_{ir}$ and $F_{vi}$ into their corresponding $Q$, $K$, and $V$, expressed as:
$\{Q_{ir},K_{ir},V_{ir}\}$ $=\mathcal{F}_{ir}^{qkv}(F_{ir}),$ (12)
$\{Q_{vi},K_{vi},V_{vi}\}$ $=\mathcal{F}_{vi}^{qkv}(F_{vi}).$
Subsequently, we swap the queries $Q$ of complementary modalities to promote spatial interaction:
$F_{f}^{ir}\!=\!\text{softmax}\Big{(}\!\frac{Q_{vi}K_{ir}}{\sqrt{d_{k}}}\!\Big{)}V_{ir},F_{f}^{vi}\!=\!\text{softmax}\Big{(}\!\frac{Q_{vi}K_{ir}}{\sqrt{d_{k}}}\!\Big{)}V_{vi},$ (13)
where $d_{k}$ is a learnable scaling factor. Finally, we concatenate $F_{f}^{ir}$ and $F_{f}^{vi}$ to get fusion features: $F_{f}=[F_{f}^{ir},F_{f}^{vi}]$.
Prompt-modulated Module. To dynamically adapt fusion feature distributions to varying degradation patterns and degrees, we propose a prompt-modulated module (PMM) for robust feature enhancement. First, sequential MLPs ($\Phi_{p}$) are utilized to derive distributional optimization parameters:
$[\gamma_{p},\beta_{p}]=\Phi_{p}(p),\quad\text{where}\ p\in\{p_{\text{vis}},p_{\text{text}}\}.$
Then, $\gamma_{p}$ and $\beta_{p}$ are applied for feature scaling and bias shifting in a residual manner:
$\hat{F}_{f}=(1+\gamma_{p})\odot F_{f}+\beta_{p},$ (14)
where $\odot$ denotes the Hadamard product, and $\hat{F}_{f}$ indicates the enhanced features incorporating degradation prompts. Image Decoder. We further deploy a series of Transformer-based decoder $\mathcal{D}_{f}$ with the self-attention mechanism to progressively reconstruct the fused image $I_{f}$:
$I_{f}=\mathcal{D}(\hat{F}_{f}).$ (15)
In particular, PMM and $\mathcal{D}_{f}$ are tightly coupled in a multi-stage process, effectively incorporating fusion features and degradation prompts to synthesize the desired image.
Table 1: Quantitative comparison results on typical fusion datasets. The best and second-best results are highlighted in Red and Purple.
MSRS LLVIP RoadScene FMB
Methods EN SD VIF Qabf EN SD VIF Qabf EN SD VIF Qabf EN SD VIF Qabf
DDFM [49] 6.431 47.815 0.844 0.643 6.914 48.556 0.693 0.517 6.994 47.094 0.775 0.595 6.426 40.597 0.495 0.442
DRMF [32] 6.268 45.117 0.669 0.550 6.901 50.736 0.786 0.626 6.231 44.221 0.728 0.527 6.842 41.816 0.578 0.372
EMMA [50] 6.747 52.753 0.886 0.605 6.366 47.065 0.743 0.547 6.959 46.749 0.698 0.664 6.788 38.174 0.542 0.436
LRRNet [12] 6.761 49.574 0.713 0.667 6.191 48.336 0.864 0.575 7.185 46.400 0.756 0.658 6.432 48.154 0.501 0.368
SegMiF [16] 7.006 57.073 0.764 0.586 7.260 45.892 0.539 0.459 6.736 48.975 0.629 0.584 6.363 47.398 0.539 0.482
Text-IF [38] 6.619 55.881 0.753 0.656 6.364 49.868 0.859 0.566 6.836 47.596 0.634 0.609 7.397 47.726 0.568 0.528
Text-Difuse [41] 6.990 56.698 0.850 0.603 7.546 55.725 0.883 0.659 6.826 50.230 0.683 0.662 6.888 49.558 0.793 0.653
ControlFusion 7.340 60.360 0.927 0.718 7.354 56.631 0.968 0.738 7.421 51.759 0.817 0.711 7.036 50.905 0.872 0.730
Table 2: Quantitative comparison results in different degraded scenarios with pre-enhancement.
VI (Blur) VI (Rain) VI (Low-light, LL) VI (Over-exposure, OE)
Methods CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD
DDFM [49] 0.141 39.421 39.047 35.411 0.191 38.836 46.285 36.376 0.156 39.495 41.782 31.759 0.143 43.167 43.440 32.099
DRMF [32] 0.128 40.739 40.968 40.722 0.174 48.164 48.565 41.174 0.143 41.428 37.947 38.287 0.190 48.334 42.582 44.256
EMMA [50] 0.131 43.472 41.744 42.553 0.138 45.824 44.916 43.378 0.158 39.674 44.827 40.857 0.180 46.731 47.616 40.242
LRRNet [12] 0.163 42.981 37.268 45.389 0.185 43.291 41.891 46.285 0.164 40.486 34.836 41.639 0.160 42.548 48.414 42.190
SegMiF [16] 0.152 43.005 43.516 44.000 0.195 40.528 49.094 44.274 0.177 44.073 48.376 44.829 0.166 49.132 38.019 38.484
Text-IF [38] 0.164 44.801 46.542 48.401 0.164 41.287 47.380 49.298 0.163 41.096 49.174 47.257 0.172 40.298 45.599 47.330
Text-Difuse [41] 0.172 44.958 47.699 46.376 0.173 39.243 50.017 47.297 0.192 46.734 50.126 49.883 0.183 39.095 49.596 50.279
ControlFusion 0.184 47.849 50.240 50.287 0.196 52.319 52.465 50.901 0.183 48.420 51.072 53.787 0.191 50.301 52.961 54.218
VI (Rain and Haze, RH) IR (Low-contrast, LC) IR (Random noise, RN) IR (Stripe noise, SN)
Methods CLIP-IQA MUSIQ TReS EN CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS EN CLIP-IQA MUSIQ TReS EN
DDFM [49] 0.158 43.119 44.095 6.897 0.172 44.490 44.545 37.452 0.186 34.059 32.807 6.029 0.209 48.479 32.377 6.399
DRMF [32] 0.171 45.524 43.693 6.230 0.208 43.982 45.561 40.315 0.192 44.617 44.085 5.744 0.187 44.714 43.650 6.354
EMMA [50] 0.169 39.092 46.046 6.517 0.154 48.724 43.113 53.861 0.172 39.666 44.466 6.184 0.153 42.802 44.239 6.382
LRRNet [12] 0.170 48.571 48.973 7.363 0.151 48.718 45.266 51.605 0.158 48.274 37.090 7.295 0.137 46.511 36.610 7.702
SegMiF [16] 0.147 46.139 45.019 7.281 0.160 44.659 51.427 44.448 0.180 42.664 39.496 6.669 0.169 49.887 39.247 6.855
Text-IF [38] 0.178 50.568 50.271 6.956 0.187 49.299 49.266 47.032 0.169 46.647 48.491 6.256 0.161 49.019 47.755 6.085
Text-Difus[41] 0.175 52.788 53.073 7.470 0.165 50.092 51.715 39.429 0.203 49.278 49.479 6.982 0.194 51.762 49.288 7.089
ControlFusion 0.189 54.287 54.465 7.891 0.196 51.986 52.846 57.827 0.189 50.711 51.668 7.724 0.200 50.097 50.264 7.619
VI (OE) and IR (LC) VI(Low-light and Noise, LN) VI (RH) and IR (RN) VI (LL) and IR (SN)
Methods CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS EN CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS EN
DDFM [49] 0.168 43.814 41.894 36.095 0.172 48.293 31.791 6.298 0.151 33.440 32.134 37.342 0.189 36.433 42.630 5.776
DRMF [32] 0.184 42.399 39.374 40.847 0.201 44.363 43.063 5.875 0.174 43.663 43.858 37.997 0.142 38.241 41.049 5.280
EMMA [50] 0.130 39.892 42.076 43.362 0.174 42.201 43.382 5.838 0.165 39.146 44.458 51.205 0.130 37.367 43.888 6.318
LRRNet [12] 0.136 47.209 42.636 46.684 0.144 46.386 35.779 7.306 0.128 47.954 36.831 49.917 0.154 38.426 35.970 7.007
SegMiF [16] 0.114 44.021 42.256 33.647 0.136 49.178 38.570 5.819 0.147 42.354 39.156 31.717 0.151 41.287 37.079 6.767
Text-IF [38] 0.174 48.808 47.998 48.848 0.217 48.100 47.510 5.204 0.158 45.821 47.626 46.543 0.140 41.429 46.220 5.525
Text-Difuse [41] 0.131 49.021 50.980 47.640 0.185 50.775 48.610 6.440 0.181 48.645 48.937 38.808 0.161 47.734 48.448 6.738
ControlFusion 0.187 50.479 50.298 50.955 0.225 49.333 49.513 7.111 0.179 50.107 51.091 55.417 0.167 50.632 48.971 7.055
4.3.2 Loss Functions
Following the typical fusion paradigm [21, 38], we introduce the intensity loss, structural similarity (SSIM) loss [46], maximum gradient loss, and color consistency loss to constrain the training of Stage II.
The intensity loss $\mathcal{L}_{int}$ aims to maximize the intensity of fusion results to ensure target prominence, defined as:
${\mathcal{L}_{int}=\frac{1}{HW}\|I_{f}-\max(I_{ir}^{hq},I_{vi}^{hq})\|_{1},}$ (16)
where $I_{ir}^{hq}$ and $I_{vi}^{hq}$ are the high-quality source images in our DDL-12 dataset. The structural similarity loss $\mathcal{L}_{ssim}$ ensures the fused image maintains structural consistency with the high-quality source images, preserving essential structural information, and is formulated as:
$\mathcal{L}_{ssim}=2-(\text{SSIM}(I_{f},I_{ir}^{hq})+\text{SSIM}(I_{f},I_{vi}^{hq})).$ (17)
The maximum gradient loss $\mathcal{L}_{grad}$ maximizes the retention of key edge information from both source images, generating fusion results with clearer textures, formulated as:
${\mathcal{L}_{grad}=\frac{1}{HW}\|\nabla I_{f}-\max(\nabla I_{ir}^{hq},\nabla I_{vi}^{hq})\|_{1},}$ (18)
where $\nabla$ denotes the Sobel operator. Moreover, the color consistency loss $\mathcal{L}_{color}$ ensures that the fusion results maintain color consistency with the visible image. We convert the image to YCbCr space and minimize the distance between the Cb and Cr channels, expressed as:
${\mathcal{L}_{color}=\frac{1}{HW}\|\mathcal{F}_{CbCr}(I_{f})-\mathcal{F}_{CbCr}(I_{vi}^{hq})\|_{1},}$ (19)
where $\mathcal{F}_{CbCr}$ denotes the transfer function of RGB to CbCr.
Finally, the total loss $\mathcal{L}_{II}$ for Stage II is the weighted sum of the aforementioned losses:
$\mathcal{L}_{II}=\alpha_{int}(p)\cdot\mathcal{L}_{int}+\alpha_{ssim}(p)\cdot\mathcal{L}_{ssim}(t)$ (20)
$+\alpha_{grad}(p)\cdot\mathcal{L}_{grad}+\alpha_{color}(p)\cdot\mathcal{L}_{color},$
where $\alpha_{int}(p)$, $\alpha_{ssim}(p)$, $\alpha_{grad}(p)$, and $\alpha_{color}(p)$ are semantically regulated hyper-parameters related to the degradation prompt $p$. In particular, we dynamically adjust the weights based on degradation characteristics (type and severity), allowing our framework to adaptively focus on both fine-grained details and holistic semantic coherence, thereby effectively addressing diverse degradations.
Figure 4: Visualization of fusion results in different degraded scenarios with pre-enhancement.
5 Experiments
5.1 Experimental Details
Implemantation Details. Our image restoration and fusion network adopts a four-tier encoder-decoder architecture with channel dimensions progressively increasing from 48 to 384, specifically configured as [48, 96, 192, 384]. During training, $224\times 224$ pixel patches are randomly cropped as inputs, using a batch size of 12 for 100 epochs. Optimization is performed using the AdamW with an initial learning rate of $1\times 10^{-3}$ and gradually reduced to $1\times 10^{-5}$ following a cosine annealing schedule. For the loss configuration, $\lambda_{1}$ and $\lambda_{2}$ are set at a $1:3$ ratio during Stage I. Additionally, $\alpha_{int}(p)$, $\alpha_{ssim}(p)$, $\alpha_{grad}(p)$, and $\alpha_{color}(p)$ are dynamically adjusted based on degradation prompts.
Experiment Configurations. We first evaluate the fusion performance of ControlFusion on four widely used datasets, i.e., MSRS [30], LLVIP [8], RoadScene [35], and FMB [16]. The test set sizes for MSRS, LLVIP, RoadScene, and FMB are 361, 50, 50, and 50, respectively. Four metrics, i.e., EN, SD, VIF, and $Q_{abf}$, are used to quantify fusion performance. We compare our method with seven SOTA fusion methods, i.e., DDFM [49], DRMF [32], EMMA [50], LRRNet [12], SegMiF [16], Text-IF [38], and Text-Difuse [41]. Furthermore, we verify the performance of ControlFusion across various degradations, including blur, rain, over-exposure, and low-light in visible images (VI), as well as low-contrast, random noise, and stripe noise in infrared images (IR). Additionally, we evaluate its effectiveness in handling composite degradation scenarios, such as rain and haze in VI, low light and noise in VI, overexposure in VI, low contrast in IR, low light in VI and stripe noise in IR, as well as rain and haze in VI combined with noise in IR. Each degradation scenario includes 100 test image pairs. We use CLIP-IQA [34], MUSIQ [10], TReS [3], EN, and SD to quantify both restoration and fusion performance. For a fair comparison, we exclusively use the visual prompts to characterize degradation types and severity in the quantitative comparison. Meanwhile, we employ several SOTA image restoration algorithms as preprocessing steps to address specific degradations. Specifically, OneRestore [6] is used for weather-related degradations, Instruct-IR [2] for illumination-related degradations, and DA-CLIP [18] and WDNN [4] for sensor-related degradations. For composite degradations, we select the best-performing method from these algorithms. All experiments are conducted on NVIDIA RTX 4090 GPUs with an Intel(R) Xeon(R) Platinum 8180 CPU (2.50 GHz) using the PyTorch framework.
5.2 Fusion Performance Comparison
Comparison without Pre-enhancement. Tab. 1 summarizes the quantitative assessment across benchmark datasets. Notably, our method attains SOTA performance in SD, VIF, and $Q_{abf}$ metrics, showcasing its exceptional ability to maintain structural integrity and edge details. The EN metric remains competitive, indicating that our fusion results encapsulate abundant information.
Comparison with Pre-enhancement. The quantitative results in Tab. 2 show that ControlFusion achieves superior CLIP-IQA, MUSIQ, and TReS scores across most degradation scenarios, demonstrating a strong capability in degradation mitigation and complementary context aggregation. Additionally, no-reference fusion metrics (SD, EN) show that ControlFusion performs on par with SOTA methods. Qualitative comparisons in Fig. 4 indicate that ControlFusion not only excels in addressing single-modal degradation but also effectively tackles more challenging single-modal and multi-modal composite degradation scenarios. In particular, when rain and haze coexist in the visible image, ControlFusion successfully perceives these degradations and eliminates their effects. In multi-modal composite degradation scenarios, it leverages degradation prompts to adjust the distribution of fusion features, achieving high-quality restoration and fusion. Extensive experiments demonstrate that ControlFusion exhibits significant advantages in complex degradation scenarios.
Figure 5: Fusion results with various level prompt.
Figure 6: Comparison of visual and textual embedding
Figure 7: Visual comparison of object detection.
5.3 Extended Experiments and Discussions
Flexible Degree Control. As shown in  Fig. 5, with degradation level prompts, our ControlFusion can adapt to varying degrees of degradation and deliver satisfactory fusion results. Moreover, fusion results generated using adjacent degradation-level prompts exhibit subtle yet perceptible differences, allowing users to obtain customized outcomes through their specific prompts.
Visual Embedding versus Textural Embedding. From Fig. 6, we can find that both visual and textual prompts assist $N_{rf}$ in restoring the scene illumination. Besides, the residual map of both is subtle, indicating that our SFVA successfully captures the degradation type and degree.
Object Detection. We further evaluate object detection performance on LLVIP to indirectly assess the fusion quality, using the re-trained YOLOv8 [27]. Qualitative and quantitative results are presented in Fig. 7 and Tab. 3. Due to effective information restoration and integration, our fusion results enable the detector to identify all pedestrians with higher confidence, achieving the highest mAP@0.5-0.95 across various confidence thresholds.
Table 3: Quantitative comparison of object detection.
Methods Prec. Recall AP@0.50 AP@0.75 mAP@0.5:0.95
DDFM [49] 0.947 0.848 0.911 0.655 0.592
DRMF [32] 0.958 0.851 0.937 0.672 0.607
EMMA [50] 0.942 0.872 0.927 0.647 0.598
LRRNet [12] 0.939 0.878 0.933 0.672 0.608
SegMiF [16] 0.965 0.896 0.931 0.690 0.603
Text-IF [38] 0.959 0.892 0.939 0.655 0.601
Text-Difuse [41] 0.961 0.885 0.941 0.656 0.606
ControlFusion 0.971 0.889 0.949 0.685 0.609
Real-World Generalization. As shown in Fig. 8, our method generalizes well to real-world scenarios, benefiting from the proposed physics-driven degraded imaging model. The first scene is captured by our multi-modal sensors with low light and noise degradation, and the second is from the FMB dataset, affected by extreme low light.
Ablation Studies. To verify the effectiveness of our key components, we conduct ablation studies across three degradation scenarios. This study involves: (I) w/o $\mathcal{L}_{color}$, (II) w/o $\mathcal{L}_{grad}$, (III) w/o $\mathcal{L}_{int}$, (IV) w/o PMM, and (V) w/o frequency branch. As shown in Fig. 9, omitting any loss or module significantly impacts the fusion quality. Specifically, without $\mathcal{L}_{color}$, color distortion occurs, while removing $\mathcal{L}_{grad}$ leads to the loss of essential texture information. Excluding $\mathcal{L}_{int}$ fails to highlight thermal targets, and removing PMM diminishes the ability to address composite degradation. Additionally, removing the frequency branch from SFVA causes visual prompt confusion. The t-SNE results further validate the importance of frequency priors. The quantitative results in Tab. 4 also confirm that each design element is crucial for enhancing fusion performance.
Figure 8: Generalization results in the real world.
Figure 9: Visualization of ablation experiment results.
Table 4: Quantitative results of the ablation studies.
VI(LL & Noise) VI(OE) and IR(LC) VI (RH) and IR(Noise)
Configs CLIP-IQA MUSIQ TReS EN CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD
I 0.132 42.424 42.839 4.788 0.166 41.983 42.841 39.917 0.147 43.619 47.932 48.935
II 0.152 45.582 45.358 5.855 0.151 43.646 42.002 39.208 0.167 47.862 45.007 44.347
III 0.154 46.571 44.495 5.013 0.155 44.286 44.561 42.068 0.156 43.007 43.816 41.544
IV 0.129 38.960 41.310 5.414 0.172 41.743 39.125 38.748 0.118 48.882 46.245 45.910
V 0.173 45.281 46.291 6.279 0.181 45.386 47.519 46.860 0.149 46.714 48.094 46.950
Ours 0.225 49.333 49.513 7.111 0.187 50.479 50.298 50.955 0.179 50.107 51.091 55.417
6 Conclusion
This work proposes a controllable framework for image restoration and fusion leveraging language-visual prompts. Initially, we develop a physics-driven degraded imaging model to bridge the domain gap between synthetic data and real-world images, providing a solid foundation for addressing composite degradations. Moreover, we devise a prompt-modulated network that adaptively adjusts the fusion feature distribution, enabling robust feature enhancement based on degradation prompts. Prompts can either come from text instructions to support user-defined control or be extracted from source images using a spatial-frequency visual adapter embedded with frequency priors, facilitating automated deployment. Extensive experiments demonstrate that the proposed method excels in handling real-world and composite degradations, showing strong robustness across various degradation levels.
Supplementary Material
A Hyper-parameters across Various Degradations
To enable the unified restoration and fusion network to adapt to various degradation types, we not only need to use degradation cues to modulate the fusion features but also customize hyper-parameters for each set of degradation types. This ensures that the model is optimized along the optimal path. Particularly, in addressing diverse degradation scenarios encountered across different data modalities, we systematically explore and iteratively refine hyper-parameter combinations through extensive empirical experimentation, ultimately identify parameter sets that demonstrate robust performance improvements. The finalized hyper-parameter selections, representing the most effective configurations derived from this comprehensive optimization process, are explicitly tabulated in Tab. 5.
Table 5: Hyper-parameters of different degradation types
Degradations $\alpha_{int}$ $\alpha_{ssim}$ $\alpha_{grad}$ $\alpha_{color}$
Blur 3 1 15 12
Rain 5 1 14 12
Haze 4 1 8 12
Low-light 8 1 10 12
Low-contrast 8 1 10 12
Stripe noise 6 0 12 12
Over-exposure 4 0 2 7
Noise (IR &VI) 6 1 10 12
Rain & Haze 5 1 12 12
Haze & Low-light 7 1 11 12
Noise & Low-light 7 1 10 12
Table 6: Details of network architecture
Num blocks Attention heads FFN factor Basic dim
Values [2,2,2,2] [1,2,4,8] 2 48
B Details of Network Architecture
This framework uses a symmetric encoder structure to process multi-modal inputs separately. Each encoder contains a multi-stage feature extraction module, and progressive feature extraction is achieved through cascaded Transformer blocks. The decoder uses a multi-level up-sampling architecture to achieve resolution recovery, and each level fuses the feature maps of the corresponding encoding stage. We employ the ViT-L/14@336px of CLIP as the text encoder to ensure semantic consistency. The specific parameter configuration of the model is shown in  Tab. 6.
Figure 10: Visualization of continuous degradation level control.
C Computational efficiency comparison
We assess the computational efficiency of ControlFusion in comparison with SOTA methods. In scenarios where the enhancement algorithm is excluded, our method does not exhibit significant competitiveness. Nevertheless, when addressing the enhancement-fusion task, our approach demonstrates superior performance. Quantitative comparisons are presented in Tab. 7.
Table 7: Comparison of model parameters, computational complexity and average testing time of different methods
Fusion Fusion with enhancement
Methods parm.(m) flops(g) time(s) parm.(m) flops(g) time(s)
DDFM 552.700 5220.000 34.500 779.590 6018.200 34.910
DRMF 5.050 121.100 0.842 170.900 4575.800 1.979
EMMA 1.516 41.540 0.037 228.450 839.230 0.440
LRRNet 0.049 14.170 0.085 226.980 811.860 0.488
SegMiF 45.040 353.700 0.147 271.970 1151.400 0.550
Text-Difuse 157.350 4872.000 31.630 157.350 4872.000 31.630
Text-IF 89.010 1518.900 0.157 89.010 1518.900 0.157
ControlFusion 123.898 327.980 0.344 123.898 327.980 0.344
D Visualization of continuous degradation level control
Although we use four discrete degradation levels to train our prompt-modulated restoration and fusion network, thanks to the powerful text parsing ability of CLIP, our model is able to achieve continuous degree control. As shown in Fig. 10, under low-light conditions, as we increase the degradation level from 1 to 10, the exposure level of the fusion results also increases. In addition, as illustrated in Fig. 10 (c), the average brightness of the three rows also increases with the degradation level prompts.