### v3

# ControlFusion: A Controllable Image Fusion Network with Language-Vision Degradation Prompts

Abstract
Current image fusion methods struggle with real-world composite degradations and lack the flexibility to accommodate user-specific needs. To address this, we propose ControlFusion, a controllable fusion network guided by language-vision prompts that adaptively mitigates composite degradations. On the one hand, we construct a degraded imaging model based on physical mechanisms, such as the Retinex theory and atmospheric scattering principle, to simulate composite degradations and provide a data foundation for addressing realistic degradations. On the other hand, we devise a prompt-modulated restoration and fusion network that dynamically enhances features according to degradation prompts, enabling adaptability to varying degradation levels. To support user-specific preferences in visual quality, a text encoder is incorporated to embed user-defined degradation types and levels as degradation prompts. Moreover, a spatial-frequency collaborative visual adapter is designed to autonomously perceive degradations from source images, thereby reducing complete reliance on user instructions. Extensive experiments demonstrate that ControlFusion outperforms SOTA fusion methods in fusion quality and degradation handling, particularly under real-world and compound degradations. The source code is publicly available at https://github.com/Linfeng-Tang/ControlFusion.
footnotetext:  Equal contribution;  * Corresponding author.
1 Introduction
Image fusion is a crucial technique in image processing. By effectively leveraging complementary information from multiple sources, the limitations associated with information collection by single-modal sensors can be significantly mitigated [41]. Among the research areas in this field, infrared-visible image fusion (IVIF) has garnered considerable attention. IVIF integrates essential thermal information from infrared (IR) images with the intricate texture details from visible (VI) images, offering a richer and more complete depiction of the scene [46]. By harmonizing diverse information and producing visually striking results, IVIF has found extensive applications in areas such as security surveillance [47], military detection [22], scene understanding [44], and assisted driving [1]etc.
Recently, IVIF has emerged as a focal point of research, leading to rapid advancements in related methods. Based on the adopted network architectures, these methods can be categorized into convolutional neural network-based [29, 51], autoencoder-based [10, 11], generative adversarial network-based [19, 15], Transformer-based [20, 45], and diffusion model-based [50, 32] methods. In addition, from a functional perspective, they can be grouped into visual-oriented [29, 51], joint registration-fusion [27, 37], semantic-driven [28, 16], and degradation-robust [31, 42, 43] schemes.
Figure 1: Comparison across real-world, composite degradation, and varying degradation levels.
Notably, although degradation-robust schemes can mitigate degradation to some extent, certain limitations exist. First, existing restoration-fusion methods often adopt simplistic strategies for training data construction, overlooking the domain gap between simulated data and realistic images, which hampers their generalizability in practical scenarios, as illustrated in Fig. 1 (I). Second, they are tailored for specific or single types of degradation, making them ineffective in handling more complex composite degradations, as illustrated in Fig. 1 (II). Finally, as shown in Fig. 1 (III), existing methods lack degradation level modeling, causing a sharp decline in performance as degradation intensifies. Moreover, they lack flexibility to adapt fusion results to diverse user preferences.
To overcome these limitations, we propose a versatile and controllable fusion model based on language-vision prompts, termed ControlFusion. On the one hand, we introduced a physics-driven degradation imaging model that differs from existing approaches by simultaneously simulating degradation processes for both visible and infrared modalities with high precision. This model not only effectively narrows the gap between synthetic data and real-world images but also provides crucial data support for addressing complex multi-modal composite degradation challenges. On the other hand, we develop a prompt-modulated image restoration and fusion network to generate high-quality fusion results. The prompt-modulated module enables our network to dynamically adjust feature distribution based on degradation characteristics (which can be specified by users), achieving robust feature enhancement. This also allows our method to respond to diverse user requirements. Furthermore, we devise a spatial-frequency visual adapter that combines frequency-domain degradation priors to directly extract degradation cues from degraded inputs, eliminating the heavy reliance on user instructions, as customizing prompts for each scene is time-consuming and labor-intensive. As shown in Fig. 1, benefiting from the aforementioned designs, our method excels in handling both real-world and composite degradation scenarios and can flexibly respond to user needs. In summary, our main contributions are as follows:
$\bullet$ We propose ControlFusion, a versatile image restoration and fusion framework that uniformly models diverse degradation types and degrees, using textual and visual prompts as a medium. Specifically, its controllability enables it to respond to user-specific customization needs.
$\bullet$ A spatial-frequency visual adapter is devised to integrate frequency characteristics and directly extract text-aligned degradation prompts from visual images, enabling automated deployment.
$\bullet$ A physics-driven imaging model is developed, integrating physical mechanisms such as the Retinex theory and atmospheric scattering principle to bridge the gap between synthetic data and real-world images, while taking into account the degradation simulation of infrared-visible dual modalities.
2 Related Work
Image Fusion. Earlier visual-oriented fusion approaches primarily concentrated on merging complementary information from multiple modalities and improving visual fidelity. The mainstream network architectures primarily include CNN-based [14, 48], AE-based [11, 49], GAN-based [19, 15], Transformers [20] and diffusion models [50]. Furthermore, several schemes including joint registration and fusion [33, 37], semantic-driven [28, 15], and degradation-robust [39, 42, 38] methods, are proposed to broaden the practical applications of image fusion. For instance, Tang et al. [30] and Liu et al. [17] developed corresponding solutions for illumination distortions and noise interference, respectively. In addition, [39] proposed a image restoration and fusion network to address various degradations. However, it fails to handle mixed degradations and struggles to generalize to real-world scenarios. In particular, it relies on tedious manual efforts to customize text prompts for each scene, hindering large-scale automated deployment.
Image Restoration with Vision-Language Models. With the advancement of deep learning toward multimodal integration, the image restoration field has progressively evolved into a new paradigm of text-driven schemes. CLIP [25] establishes visual-textual semantic representation alignment through dual-stream Transformer encoders pre-trained on 400 million image-text pairs. This framework lays the foundation for the widespread adoption of Prompt Engineering in computer vision. Recent studies integrate CLIP encoders with textual user instructions, proposing generalized restoration frameworks including PromptIR [23], AutoDIR [8], and InstructIR [2], which effectively handle diverse degradation types. Moreover, SPIRE [24] further introduces fine-grained textual restoration cues by quantifying degradation severity levels and supplementing semantic information for precision restoration. To eliminate reliance on manual guidance, Luo et al. [18] developed DA-CLIP by fine-tuning CLIP on mixed-degradation datasets, enabling degradation-aware embeddings to facilitate distortion handling. However, existing unified restoration methods are primarily designed for natural images, which exhibit notable limitations when processing multimodal images (e.g., infrared-visible).
3 Physics-driven Degraded Imaging Model
Due to the differences in the imaging mechanisms of infrared and visible, the types of degradation they face also vary. To tackle the complex and variable degradation challenges, we propose a physics-driven imaging model for infrared and visible images aimed at reducing the domain gap between simulated data and real-world imagery. Infrared (IR) images usually suffer from sensor-related interference, such as stripe noise and low contrast. While Visible (VI) images are typically degraded by illumination conditions (low light, over-exposure), weather (rain, haze), and sensor-related issues (noise, blur). Given a clear image $I_{m}(m\in\{ir,vi\})$, the proposed imaging model can be mathematically represented as:
$D_{m}=\mathcal{P}_{s}\left(\mathcal{P}_{w}\left(\mathcal{P}_{i}\left(I_{m}\right)\right)\right),$ (1)
where $D_{m}$ is the corresponding degraded image, and $\mathcal{P}_{i}$, $\mathcal{P}_{w}$, and $\mathcal{P}_{s}$ represent illumination-, weather-, and sensor-related distortions, respectively.
Sensor-related Distortions. Sensor-related distortions encompass various types of noise, motion blur, and contrast degradation. Contrast degradation and stripe noise of infrared images can be modeled as:
$D^{s}_{ir}=\mathcal{P}_{s}(I_{ir})=\alpha\cdot I_{ir}+\mathbf{1}_{H}\mathbf{n}^{\top},$ (2)
where $\alpha$ is a constant less than 1 for contrast reduction, $\mathbf{1}_{H}\in\mathbb{R}^{H}$ is an all-ones column vector and $\mathbf{n}\in\mathbb{R}^{W}$ represents the column-wise Gaussian noise vector sampled from $\mathcal{N}(0,\epsilon^{2})$ with $\epsilon\in[1,15]$. Moreover, Gaussian noise and motion blur in source images can be modeled as:
$D^{s}_{m}=\mathcal{P}_{s}(I_{m})=I_{m}\ast K(N,\theta)+\mathcal{N}(0,\sigma^{2}),$ (3)
where $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise with $\sigma\in[5,20]$. * denotes convolution with a blur kernel $K(N,\theta)=\frac{1}{N}R_{\theta}(\delta_{c}\otimes\mathbf{1}_{N})$ constructed by rotating an impulse $\delta_{c}\otimes\mathbf{1}_{N}$ with random angle $R_{\theta}$ ($\theta\in[10,80]$), using $\frac{1}{N}$ to normalize the kernel energy. Here, $N\in[3,12]$ controls the blur level.
Illumination-related Distortions. Following the theoretical framework of Retinex, the formation of illumination-degraded images $D^{i}_{vi}$ is mathematically modeled as:
${D^{i}_{vi}=\mathcal{P}_{i}(I_{vi})=\frac{I_{vi}}{L}\cdot L^{\gamma},}$ (4)
where $L$ is the illumination map estimated by LIME [5]. $\gamma\in[0.5,3]$ controls the illumination level.
Weather-related Distortions. According to the methodologies in [21, 12], we employ the following formula to simulate weather-related degradations (i.e. rain and haze):
$D^{w}_{vi}=\mathcal{P}_{w}(I_{vi})=I_{vi}\cdot t+A(1-t)+R,$ (5)
where $t$ and $A$ denote the transmission map and atmospheric light, respectively. And $t$ is defined by the exponential decay of light, expressed as $t=e^{-\beta d(x)}$, where the haze density coefficient $\beta\in[0.5,2.0]$. $d(x)$ refers to the scene depth, estimated by DepthAnything-V2 [40]. Besides, $A$ is constrained within $[0.3,0.9]$ for realistic simulation.
Based on Eq. (1), we construct a multi-modal composite Degradation Dataset with four Levels (DDL-12), comprising 12 distinct degradations. We selecte $2,050$ high-quality clear images from RoadScene [36], LLVIP [7], and MSRS [29] datasets. Subsequently, the degraded imaging model is employed to synthesize degraded images with 12 types and 4 levels of degradation. Finally, the DDL-12 dataset contains approximately $48,000$ training image pairs and $4,800$ test image pairs.
Figure 2: The overall framework of our controllable image fusion network.
4 Methodology
4.1 Problem Formulation
Given two source images $I_{ir}\in\mathbb{R}^{H\times W\times 1}$ and $I_{vi}\in\mathbb{R}^{H\times W\times 3}$, the typical fusion paradigm employs a network $\mathcal{N}_{f}$ to synthesize the fused image $I_{f}$, expressed as: $I_{f}=\mathcal{N}_{f}(I_{ir},I_{vi})$. However, in complex scenarios, source images usually suffer from degradation interference, thus the advanced fusion paradigm must take both image restoration and fusion into account. While the concatenation approach is straightforward, it does not model recovery and fusion as an end-to-end optimization process, yielding suboptimal outcomes. As illustrated in Fig. 2, we propose a controllable paradigm that couples restoration and fusion with degradation prompts, termed ControlFusion, to overcome this limitation. On the one hand, the coupled approach enhances task synergy. On the other hand, leveraging degradation prompts to modulate the restoration and fusion process enables the network to adapt to diverse degradation distributions while meeting user-specific customization requirements. The proposed restoration and fusion paradigm can be defined as: $I_{f}=\mathcal{N}_{rf}(I_{ir},I_{vi},p~|~\Omega)$, where $\mathcal{N}_{rf}$ is a restoration and fusion network, $P$ denotes degradation prompts, and $\Omega$ indicates the degradation set. Our ControlFusion employs a two-stage optimization protocol. Stage I aligns textual prompts with visual embedding, while Stage II optimizes the overall framework.
4.2 Stage I: Textual-Visual Prompts Alignment
Spatial-Frequency Collaborative Visual Adapter. To achieve unified multimodal degradation representation, we first employ text semantics for explicit degradation modeling. Using a pre-trained CLIP [25] text encoder $\mathcal{E}_{t}$, the user instruction $T_{instr}$ is converted into text embedding: $p_{text}=\mathcal{E}_{t}(T_{instr})$. However, text-dependent models limit deployment flexibility. We therefore develop a spatial-frequency collaborative visual adapter (SFVA) to extract visual embeddings directly from images while maintaining semantic alignment with $p_{text}$.
As shown in Fig. 3, degraded images exhibit distinct spectral characteristics, indicating frequency features contain rich degradation priors. Our SFVA contains dual branches: In the frequency branch, Fast Fourier Transform (FFT) and CNN extract frequency features:
$F_{fre}^{m}$ $=\textstyle{\sum_{x=0}^{W-1}\sum_{y=0}^{H-1}D_{m}(x,y)e^{-j2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)},}~F_{fre}=\text{Linear}\left(\text{Conv}\left([F_{fre}^{ir},F_{fre}^{vi}]\right)\right),$ (6)
where $\cdot$$\cdot$] denotes channel-wise concatenation. The spatial branch employs cropping/downsampling augmentation and CNN to extract spatial features $F_{spa}$.
The fused visual embedding $p_{vis}$ is obtained through concatenation and linear projection of $F_{fre}$ and $F_{spa}$. To ensure semantic consistency between visual/text embeddings, we apply MSE loss $\mathcal{L}_{mse}$ and cosine similarity loss $\mathcal{L}_{cos}$:
$\mathcal{L}_{I}=\lambda_{1}\underbrace{\|p_{vis}-p_{text}\|^{2}}_{\mathcal{L}_{\text{mse}}}+\lambda_{2}\underbrace{(1-\frac{p_{vis}\cdot p_{text}}{\|p_{vis}\|\|p_{text}\|})}_{\mathcal{L}_{\text{cos}}},$ (7)
where $\lambda_{1}$ and $\lambda_{2}$ balance loss components. This design enables automatic degradation-aware adaptation while preserving text-aligned semantics.
Figure 3: The visualization of various types of degradation in the spatial and frequency domains.
4.3 Stage II: Prompt-modulated Restoration and Fusion
4.3.1 Network Architectures
[…]
Furthermore, to achieve comprehensive cross-modal feature integration, we design the intra-domain fusion unit, where the cross-attention (CR-ATT) mechanism is employed to facilitate interaction between features across modalities. Specifically, the linear transformation functions $\mathcal{F}_{ir}^{qkv}$ and $\mathcal{F}_{vi}^{qkv}$ project $F_{ir}$ and $F_{vi}$ into their corresponding $Q$, $K$, and $V$, expressed as:
$\{Q_{ir},K_{ir},V_{ir}\}$ $=\mathcal{F}_{ir}^{qkv}(F_{ir}),\{Q_{vi},K_{vi},V_{vi}\}$ $=\mathcal{F}_{vi}^{qkv}(F_{vi}).$ (8)
Subsequently, we swap the queries $Q$ of complementary modalities to promote spatial interaction:
$F_{f}^{ir}\!=\!\text{softmax}\Big(\!\frac{Q_{vi}K_{ir}}{\sqrt{d_{k}}}\!\Big)V_{ir},F_{f}^{vi}\!=\!\text{softmax}\Big(\!\frac{Q_{ir}K_{vi}}{\sqrt{d_{k}}}\!\Big)V_{vi},$ (9)
where $d_{k}$ is scaling factor. We concatenate $F_{f}^{ir}$ and $F_{f}^{vi}$ to get fusion features: $F_{f}=[F_{f}^{ir},F_{f}^{vi}]$.
[…]
4.3.2 Loss Functions
Following the typical fusion paradigm [20], we introduce the intensity loss, structural similarity (SSIM) loss, maximum gradient loss, and color consistency loss to constrain the training of Stage II.
The intensity loss $\mathcal{L}_{int}$ maximizes the target prominence of fusion results, defined as:
${\mathcal{L}_{int}=\frac{1}{HW}\|I_{f}-\max(I_{ir}^{hq},I_{vi}^{hq})\|_{1},}$ (10)
where $I_{ir}^{hq}$ and $I_{vi}^{hq}$ are the high-quality source images.
The structural similarity loss $\mathcal{L}_{ssim}$ ensures the fused image maintains structural consistency with the high-quality source images, preserving essential structural information, and is formulated as:
$\mathcal{L}_{ssim}=2-(\text{SSIM}(I_{f},I_{ir}^{hq})+\text{SSIM}(I_{f},I_{vi}^{hq})).$ (11)
The maximum gradient loss $\mathcal{L}_{grad}$ maximizes the retention of key edge information from both source images, generating fusion results with clearer textures, formulated as:
${\mathcal{L}_{grad}=\frac{1}{HW}\|\nabla I_{f}-\max(\nabla I_{ir}^{hq},\nabla I_{vi}^{hq})\|_{1},}$ (12)
where $\nabla$ denotes the Sobel operator. Moreover, the color consistency loss $\mathcal{L}_{color}$ ensures that the fusion results maintain color consistency with the visible image. We convert the image to YCbCr space and minimize the distance between the Cb and Cr channels, expressed as:
${\mathcal{L}_{color}=\frac{1}{HW}\|\mathcal{F}_{CbCr}(I_{f})-\mathcal{F}_{CbCr}(I_{vi}^{hq})\|_{1},}$ (13)
where $\mathcal{F}_{CbCr}$ denotes the transfer function of RGB to CbCr.
Finally, the total loss $\mathcal{L}_{II}$ for Stage II is the weighted sum of the aforementioned losses:
$\mathcal{L}_{II}=\alpha_{int}\cdot\mathcal{L}_{int}+\alpha_{ssim}\cdot\mathcal{L}_{ssim}+\alpha_{grad}\cdot\mathcal{L}_{grad}+\alpha_{color}\cdot\mathcal{L}_{color},$ (14)
where $\alpha_{int}$, $\alpha_{ssim}$, $\alpha_{grad}$, and $\alpha_{color}$ are hyper-parameters.
5 Experiments
5.1 Implementation and Experimental Configurations
Our image restoration and fusion network is built on a four-stage encoder–decoder architecture, with channel dimensions increasing from 48 to 384, specifically configured as $[48,96,192,384]$. The model is trained on the proposed DDL-12 dataset. During training, $224\times 224$ patches are randomly cropped as inputs, with a batch size of 12 over 100 epochs. Optimization is performed using AdamW, starting with a learning rate of $1\times 10^{-3}$ and decayed to $1\times 10^{-5}$ via a cosine annealing schedule. For the loss configuration, $\lambda_{1}$ and $\lambda_{2}$ are set with a weight ratio of $1:3$, and $\alpha_{int}$, $\alpha_{ssim}$, $\alpha_{grad}$, and $\alpha_{color}$ are assigned values of $8:1:10:12$, respectively.
We introduce text prompts to enable the unified network to effectively handle diverse and complex degradations. Each prompt specifies the affected modality, the degradation type, and its severity, enabling user-controllable flexibility. For example, a typical prompt for a single degradation is: We are performing infrared and visible image fusion, where the modality suffers from a grade-severity degradation type. For composite degradations, we extend this template to specify multiple modality–degradation pairs, e.g., We are performing infrared and visible image fusion. Please handle a grade-severity-A degradation type-A in the modality-A, and a grade-severity-B degradation type-B in the modality-B. Details of the prompt construction paradigm are provided in the Appendix.
We compare our method with seven SOTA fusion methods, i.e., DDFM [50], DRMF [31], EMMA [51], LRRNet [11], SegMiF [16], Text-IF [39], and Text-DiFuse [42]. Firstly, we evaluate the fusion performance on four widely used datasets, i.e., MSRS [29], LLVIP [7], RoadScene [35], and FMB [16]. The test set sizes for MSRS, LLVIP, RoadScene, and FMB are 361, 50, 50, and 50, respectively. Four metrics, i.e., EN, SD, VIF, and $Q_{abf}$, are used to quantify fusion performance. Additionally, we evaluate the restoration and fusion performance across various degradations, including low-contrast, random noise, and stripe noise in infrared images (IR), as well as blur, rain, over-exposure, and low light in visible images (VI). Additionally, we evaluate its effectiveness in handling composite degradation scenarios. Each degradation scenario includes 100 image pairs from DDL-12 dataset. We use CLIP-IQA [34], MUSIQ [9], TReS [3], EN, and SD to quantify both restoration and fusion performance. All experiments are conducted on NVIDIA RTX 4090 GPUs with an Intel(R) Xeon(R) Platinum 8180 CPU (2.50 GHz) using the PyTorch framework.
Table 1: Quantitative comparison results on typical fusion datasets. The best and second-best results are highlighted in Red and Purple.
MSRS LLVIP RoadScene FMB
Methods EN SD VIF Qabf EN SD VIF Qabf EN SD VIF Qabf EN SD VIF Qabf
DDFM 6.431 47.815 0.844 0.643 6.914 48.556 0.693 0.517 6.994 47.094 0.775 0.595 6.426 40.597 0.495 0.442
DRMF 6.268 45.117 0.669 0.550 6.901 50.736 0.786 0.626 6.231 44.221 0.728 0.527 6.842 41.816 0.578 0.372
EMMA 6.747 52.753 0.886 0.605 6.366 47.065 0.743 0.547 6.959 46.749 0.698 0.664 6.788 38.174 0.542 0.436
LRRNet 6.761 49.574 0.713 0.667 6.191 48.336 0.864 0.575 7.185 46.400 0.756 0.658 6.432 48.154 0.501 0.368
SegMiF 7.006 57.073 0.764 0.586 7.260 45.892 0.539 0.459 6.736 48.975 0.629 0.584 6.363 47.398 0.539 0.482
Text-IF 6.619 55.881 0.753 0.656 6.364 49.868 0.859 0.566 6.836 47.596 0.634 0.609 7.397 47.726 0.568 0.528
Text-DiFuse 6.990 56.698 0.850 0.603 7.546 55.725 0.883 0.659 6.826 50.230 0.683 0.662 6.888 49.558 0.793 0.653
ControlFusion 7.340 60.360 0.927 0.718 7.354 56.631 0.968 0.738 7.421 51.759 0.817 0.711 7.036 50.905 0.872 0.730
Table 2: Quantitative comparison results under different degradation scenarios with pre-enhancement.
VI (Blur) VI (Rain) VI (Low light, LL) VI (Over-exposure, OE)
Methods CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD
DDFM 0.141 39.421 39.047 35.411 0.191 38.836 46.285 36.376 0.156 39.495 41.782 31.759 0.143 43.167 43.440 32.099
DRMF 0.128 40.739 40.968 40.722 0.174 48.164 48.565 41.174 0.143 41.428 37.947 38.287 0.190 48.334 42.582 44.256
EMMA 0.131 43.472 41.744 42.553 0.138 45.824 44.916 43.378 0.158 39.674 44.827 40.857 0.180 46.731 47.616 40.242
LRRNet 0.163 42.981 37.268 45.389 0.185 43.291 41.891 46.285 0.164 40.486 34.836 41.639 0.160 42.548 48.414 42.190
SegMiF 0.152 43.005 43.516 44.000 0.195 40.528 49.094 44.274 0.177 44.073 48.376 44.829 0.166 49.132 38.019 38.484
Text-IF 0.164 44.801 46.542 48.401 0.164 41.287 47.380 49.298 0.163 41.096 49.174 47.257 0.172 40.298 45.599 47.330
Text-DiFuse 0.172 44.958 47.699 46.376 0.173 39.243 50.017 47.297 0.192 46.734 50.126 49.883 0.183 39.095 49.596 50.279
ControlFusion 0.184 47.849 50.240 50.287 0.196 52.319 52.465 50.901 0.183 48.420 51.072 53.787 0.191 50.301 52.961 54.218
VI (Rain and Haze, RH) IR (Low-contrast, LC) IR (Random noise, RN) IR (Stripe noise, SN)
Methods CLIP-IQA MUSIQ TReS EN CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS EN CLIP-IQA MUSIQ TReS EN
DDFM 0.158 43.119 44.095 6.897 0.172 44.490 44.545 37.452 0.186 34.059 32.807 6.029 0.209 48.479 32.377 6.399
DRMF 0.171 45.524 43.693 6.230 0.208 43.982 45.561 40.315 0.192 44.617 44.085 5.744 0.187 44.714 43.650 6.354
EMMA 0.169 39.092 46.046 6.517 0.154 48.724 43.113 53.861 0.172 39.666 44.466 6.184 0.153 42.802 44.239 6.382
LRRNet 0.170 48.571 48.973 7.363 0.151 48.718 45.266 51.605 0.158 48.274 37.090 7.295 0.137 46.511 36.610 7.702
SegMiF 0.147 46.139 45.019 7.281 0.160 44.659 51.427 44.448 0.180 42.664 39.496 6.669 0.169 49.887 39.247 6.855
Text-IF 0.178 50.568 50.271 6.956 0.187 49.299 49.266 47.032 0.169 46.647 48.491 6.256 0.161 49.019 47.755 6.085
Text-DiFuse 0.175 52.788 53.073 7.470 0.165 50.092 51.715 39.429 0.203 49.278 49.479 6.982 0.194 51.762 49.288 7.089
ControlFusion 0.189 54.287 54.465 7.891 0.196 51.986 52.846 57.827 0.189 50.711 51.668 7.724 0.200 50.097 50.264 7.619
VI (OE) and IR (LC) VI(Low light and Noise, LN) VI (RH) and IR (RN) VI (LL) and IR (SN)
Methods CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS EN CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS EN
DDFM 0.168 43.814 41.894 36.095 0.172 48.293 31.791 6.298 0.151 33.440 32.134 37.342 0.189 36.433 42.630 5.776
DRMF 0.184 42.399 39.374 40.847 0.201 44.363 43.063 5.875 0.174 43.663 43.858 37.997 0.142 38.241 41.049 5.280
EMMA 0.130 39.892 42.076 43.362 0.174 42.201 43.382 5.838 0.165 39.146 44.458 51.205 0.130 37.367 43.888 6.318
LRRNet 0.136 47.209 42.636 46.684 0.144 46.386 35.779 7.306 0.128 47.954 36.831 49.917 0.154 38.426 35.970 7.007
SegMiF 0.114 44.021 42.256 33.647 0.136 49.178 38.570 5.819 0.147 42.354 39.156 31.717 0.151 41.287 37.079 6.767
Text-IF 0.174 48.808 47.998 48.848 0.217 48.100 47.510 5.204 0.158 45.821 47.626 46.543 0.140 41.429 46.220 5.525
Text-DiFuse 0.131 49.021 50.980 47.640 0.185 50.775 48.610 6.440 0.181 48.645 48.937 38.808 0.161 47.734 48.448 6.738
ControlFusion 0.187 50.479 50.298 50.955 0.225 49.333 49.513 7.111 0.179 50.107 51.091 55.417 0.167 50.632 48.971 7.055
5.2 Fusion Performance Comparison
Comparison without Pre-enhancement. Table 1 summarizes the quantitative assessment across benchmark datasets. Notably, our method attains SOTA performance in SD, VIF, and $Q_{abf}$ metrics, showcasing its exceptional ability to maintain structural integrity and edge details. The EN metric remains competitive, indicating that our fusion results encapsulate abundant information.
Comparison with Pre-enhancement. For a fair comparison, we exclusively use the visual prompts to characterize degradation in the quantitative comparison. Meanwhile, we employ several SOTA image restoration algorithms as preprocessing steps to address specific degradations. Specifically, OneRestore [6] is used for weather-related degradations, Instruct-IR [2] for illumination-related degradations, and DA-CLIP [18] and WDNN [4] for sensor-related degradations. For composite degradations, we select the best-performing method from these algorithms.
The quantitative results in Tab. 2 show that ControlFusion achieves superior CLIP-IQA, MUSIQ, and TReS scores across most degradation scenarios, demonstrating a strong capability in degradation mitigation and complementary context aggregation. Additionally, no-reference fusion metrics (SD, EN) show that ControlFusion performs on par with SOTA image fusion methods. Qualitative comparisons in Fig. 4 indicate that ControlFusion not only excels in addressing single-modal degradation but also effectively tackles more challenging single- and multi-modal composite degradations. In particular, when rain and haze coexist in the visible image, ControlFusion successfully perceives these degradations and eliminates their effects. In multi-modal composite degradation scenarios, it leverages degradation prompts to adjust the distribution of fusion features, achieving high-quality restoration and fusion. Extensive experiments demonstrate that ControlFusion exhibits significant advantages in complex degradation scenarios.
Figure 4: Visualization of fusion results under different degradation scenarios.
Table 3: Quantitative comparison on MSRS nighttime scenes and AWMM-100k with real-world weather degradations.
MSRS (Nighttime) AWMM-100k (Weather Degradations)
Methods CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD
DDFM 0.104 25.034 19.404 33.935 0.227 41.919 53.858 36.010
DRMF 0.102 27.518 16.052 46.718 0.232 50.054 56.671 39.735
EMMA 0.087 26.713 18.511 42.659 0.253 47.161 49.722 40.398
LRRNet 0.098 27.042 16.909 42.692 0.192 46.751 50.707 37.190
SegMiF 0.093 27.405 21.079 44.315 0.257 50.712 54.310 33.311
Text-IF 0.101 28.155 20.640 49.752 0.213 45.633 51.747 32.086
Text-DiFuse 0.116 28.620 18.967 43.728 0.233 46.901 57.708 41.530
ControlFusion 0.141 28.457 21.607 52.933 0.280 57.513 60.302 44.244
5.3 Extended Experiments
Figure 5: Generalization results under real-world degradation scenarios.
Real-world Generalization. As shown in Tab. 3, we report comparative results on nighttime scenes from MSRS and the real weather-degraded benchmark AWMM-100k [13]. Our method achieves the best performance on CLIP-IQA, TRes, and SD, while slightly trailing Text-DiFuse on MUSIQ . On AWMM-100k, our approach consistently outperforms competing methods across all evaluation metrics under real-world weather degradation conditions.
In addition to quantitative comparisons, Fig. 5 presents visual examples demonstrating our method’s strong generalization ability in removing composite degradations in real scenes. The top image shows typical data collected by our multimodal sensors under challenging low-light and noisy conditions, while the bottom image illustrates the robustness of our method in handling extremely low-light scenarios. Additional qualitative results are provided in the Appendix.
Figure 6: Fusion results with various level prompts.
Flexible Degree Control. We employ numerical values to quantify the severity of degradation in design, where higher numbers indicate more severe degradation, enabling precise continuous control. It should be noted that while only four anchor points (1, 4, 7, and 10) were used during training, the model demonstrates remarkable generalization capability to intermediate degrees during testing, owing to the inherent encoding characteristics of CLIP. As shown in Fig. 6, with degradation level prompts, our ControlFusion can adapt to varying degrees of degradation and deliver satisfactory fusion results. Moreover, fusion results generated using adjacent degradation-level prompts exhibit subtle yet perceptible differences, allowing users to obtain customized outcomes through their specific prompts. More visualization results are presented in the Appendix.
Table 4: Computational efficiency of image restoration and fusion algorithms on $640\times 480$ resolution.
Image Restoration Image Fusion
Methods Parm.(M) Flops(G) Time(s) Methods Parm.(M) Flops(G) Time(s) Methods Parm.(M) Flops(G) Time(s)
InstructIR 15.94 151.54 0.038 DDFM 552.70 5220.50 34.511 SegMiF 45.64 707.39 0.371
DA-CLIP 295.19 1214.55 13.578 DRMF 5.05 121.10 0.842 Text-IF 152.44 1518.90 0.157
WDNN 0.01 1.11 0.004 EMMA 1.52 41.54 0.048 Text-DiFuse 157.35 4872.00 12.903
OneRestore 18.12 114.56 0.024 LRRNet 0.05 14.17 0.096 ControlFusion 182.54 1622.00 0.183
Computational Efficiency. Table 4 details the computational costs for various methods. While standalone fusion models such as EMMA, LRRNet, and SegMiF, which are marked with , appear more lightweight, they necessitate a separate, computationally intensive restoration stage to form a practical pipeline. When the overhead of this two-stage process is factored in, the efficiency advantages of our unified model becomes evident. Moreover, our model’s efficiency remains highly competitive with other joint restoration-fusion approaches like DRMF, Text-DiFuse, and Text-IF.
Table 5: Quantitative comparison of object detection.
Methods Prec. Recall AP@0.50 AP@0.75 mAP@0.5:0.95
DDFM 0.947 0.848 0.911 0.655 0.592
DRMF 0.958 0.851 0.937 0.672 0.607
EMMA 0.942 0.872 0.927 0.647 0.598
LRRNet 0.939 0.878 0.933 0.672 0.608
SegMiF 0.965 0.896 0.931 0.690 0.603
Text-IF 0.959 0.892 0.939 0.655 0.601
Text-DiFuse 0.961 0.885 0.941 0.656 0.606
ControlFusion 0.971 0.889 0.949 0.685 0.609
Object Detection. We evaluate object detection performance on LLVIP to assess the fusion quality, using the re-trained YOLOv8 [26]. Quantitative results are presented in Tab. 5. Our results enable the detector to identify all pedestrians with higher confidence, achieving the highest mAP@0.5-0.95. Visualizations are in the Appendix.
Figure 7: Visual results of ablation studies under degradation scenarios.
Table 6: Quantitative results of the ablation studies.
VI(LL & Noise) VI(OE) and IR(LC) VI (RH) and IR(Noise)
Configs CLIP-IQA MUSIQ TReS EN CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD
I 0.132 42.424 42.839 4.788 0.166 41.983 42.841 39.917 0.147 43.619 47.932 48.935
II 0.152 45.582 45.358 5.855 0.151 43.646 42.002 39.208 0.167 47.862 45.007 44.347
III 0.154 46.571 44.495 5.013 0.155 44.286 44.561 42.068 0.156 43.007 43.816 41.544
IV 0.129 38.960 41.310 5.414 0.172 41.743 39.125 38.748 0.118 48.882 46.245 45.910
V 0.173 45.281 46.291 6.279 0.181 45.386 47.519 46.860 0.149 46.714 48.094 46.950
Ours 0.225 49.333 49.513 7.111 0.187 50.479 50.298 50.955 0.179 50.107 51.091 55.417
Ablation Studies. To verify the effectiveness of components, we conduct detailed ablation studies, involves: (I) w/o $\mathcal{L}_{color}$, (II) w/o $\mathcal{L}_{grad}$, (III) w/o $\mathcal{L}_{int}$, (IV) w/o PMM, and (V) w/o frequency branch. As shown in Fig. 7, removing any loss or module significantly impacts the fusion quality. Specifically, without $\mathcal{L}_{color}$, color distortion occurs, while removing $\mathcal{L}_{grad}$ leads to the loss of essential texture information. Excluding $\mathcal{L}_{int}$ fails to highlight thermal targets, and removing PMM diminishes the ability to address composite degradation. Additionally, removing the frequency branch from SFVA causes visual prompt confusion. The t-SNE results further validate the importance of frequency priors. The quantitative results in Tab. 6 also confirm that each design element is crucial for enhancing fusion performance.
Discussions and Limitations. To mitigate the gap between simulated and real-world data, our method constructs training samples using the degradation imaging model. An alternative strategy is test-time adaptation, in which part of the model is fine-tuned on the test set to better accommodate new data distributions. It should be noted that the degradation imaging model is specifically designed for infrared and visible image fusion and is difficult to generalize to other fusion tasks, such as medical image fusion and multi-focus image fusion.
6 Conclusion
This work proposes a controllable framework for image restoration and fusion leveraging language-visual prompts. Initially, we develop a physics-driven degraded imaging model to bridge the domain gap between synthetic data and real-world images, providing a solid foundation for addressing composite degradations. Moreover, we devise a prompt-modulated network that adaptively adjusts the fusion feature distribution, enabling robust feature enhancement based on degradation prompts. Prompts can either come from text instructions to support user-defined control or be extracted from source images using a spatial-frequency visual adapter embedded with frequency priors, facilitating automated deployment. Extensive experiments demonstrate that our method excels in handling real-world and composite degradations, showing strong robustness across various degradation levels.
7 Acknowledgement
This work was supported by NSFC (62276192).
Appendix A Appendix
A.1 Prompt Design Details
We provide additional details of the prompt construction paradigm to complement the main text.
Prompt components. Each prompt specifies three key components: (i) the affected modality (infrared or visible), (ii) the degradation type (e.g., rain, haze, low-light, overexposure, random noise, stripe noise, low-contrast, etc.), and (iii) the severity level (an integer from 1 to 10).
Severity levels. To characterize degradation severity, we define a scale from 1 to 10, with representative anchor points as follows:
  • Level 1: barely perceptible degradation,
  • Level 4: degradation begins to interfere with human scene understanding,
  • Level 7: degradation severely hinders human perception of the scene,
  • Level 10: most useful information is completely obscured.
Intermediate levels can be interpolated between these anchor points.
Prompt templates. A typical template for a single degradation is structured as: We are performing infrared and visible image fusion, where the modality suffers from a grade-severity degradation type.
For composite degradations, two extensions are designed: 1) We are performing infrared and visible image fusion. Please handle a grade-severity-A degradation type-A in the modality-A, and a grade-severity-B degradation type-B in the modality-B; 2) We are performing infrared and visible image fusion. Please address level-severity degradation type-A and degradation type-B in the modality.
These templates flexibly represent both intra- and inter-modal combinations. For instance, an instantiation could be: We are performing infrared and visible image fusion. Please handle a grade-6 low-light in the visible modality, and a grade-8 stripe noise in the infrared modality.
Linguistic diversity. To enhance robustness, we curated over 100 unique textual formulations for each task and its associated parameters. This prompt rephrasing strategy augments linguistic variability and improves the model’s generalization to diverse prompt styles during inference.
Additionally, when users are uncertain about severity, the proposed SFVA module can automatically perceive both degradation type and severity from the input. This design balances user-controllable flexibility with automated practicality.
A.2 Visual Embedding versus Textual Embedding
From Fig. 8, we can find that both visual and textual prompts assist $N_{rf}$ in restoring the scene illumination. Besides, the residual plots between the two are not obvious, which indicates that our SFVA successfully captures the type and extent of degradation.This capability is crucial for automatic deployment, as it allows the model to adaptively adjust its restoration strategy with fewer manual lintervention or hyperparameters tuning for different degradation scenarios.
Figure 8: Comparison of visual and textual embedding.
A.3 Flexible Degree Control
In order to verify the effectiveness of the proposed flexible degree control, we perform different degrees of prompts on a pair of low-light images. The results are shown in Fig. 9 (b). We splice all the results to obtain images with continuously changing brightness in different regions, and count the brightness of the same region of these images. The results are shown in Fig. 9 (c). This shows that our degree control achieves fine-grained regulation.
Figure 9: Fusion results with various level prompts.
A.4 Real-world Generalization
We conduct comprehensive experiments on the FMB dataset under various composite degradation scenarios. Notably, these test data are collected from real-world environments and represent degradation patterns not encountered in the training set. During evaluation, we rely exclusively on visual prompts automatically extracted from the input images, demonstrating our method’s ability to handle unseen degradation conditions without manual intervention, as shown in Fig. 10.
Figure 10: Real-world composite degradation scenario testing.
A.5 Effect of Severity Level Granularity
To ensure our model learns to discern fine-grained degradation severity, we train it on a dense set of four anchor degradation levels (1&4&7&10) from the DDL-12 dataset. This design is supported by a comparative analysis against models trained on sparser subsets of only two levels, specifically the moderate pair (4&7) or the extreme pair (1&10). As shown in Tab. 7, these sparser training regimes lead to a notable drop in generalization performance on real-world scenarios, particularly when only the extreme levels (1&10) are used. These results indicate the importance of dense sampling of training levels for robust generalization.
Table 7: Comparison of models trained with varying degradation levels on real-world datasets.
LLVIP FMB
Level numbers CLIP-IQA MUSIQ TReS SD CLIP-IQA MUSIQ TReS SD
2 (1&10) 0.320 53.498 61.889 54.665 0.192 52.839 63.115 50.122
2 (4&7) 0.332 55.364 64.382 55.742 0.207 52.771 63.328 50.429
4 (1&4&7&10) 0.347 55.890 65.014 56.657 0.208 53.034 63.704 51.707
Figure 11: Visualization results of target detection.
A.6 Object Detection
As illustrated in Fig. 11, our fusion method achieves superior object detection performance with consistently high confidence scores. Notably, the detector successfully identifies all pedestrians in challenging scenarios, including heavily occluded cases where other methods fail. These results quantitatively validate the effectiveness of our approach in preserving and enhancing critical visual information through the fusion process, particularly for small and obscured targets. The improved detection performance further confirms that our method maintains both thermal signatures and texture details essential for downstream vision tasks.
A.7 Model Complexity Analysis
As demonstrated in Fig. 7 and Tab. 6, both CLIP and SFVA (i.e., visual–text prompt alignment) yield substantial improvements in restoration and fusion performance. We analyze the computational cost of these components in Tab. 8. Although CLIP accounts for a significant number of parameters (102M, 55.9% of the total), they are frozen during training and thus introduce no training overhead. In terms of inference cost, CLIP and SFVA are responsible for a modest 0.81% and 7.5% of the total FLOPs, respectively, and their combined runtime constitutes a small fraction of the total. Overall, both components offer a compelling trade-off, delivering notable performance gains for a moderate computational expense.
Table 8: Overhead of individual components. Percentages indicate the relative share in the full model.
Components Parm. (M) FLOPs (G) Time (s)
CLIP-L/14@336px 102.00 (55.88%) 13.088 (0.81%) 0.0120 (6.55%)
SFVA 15.27 (8.37%) 122.32 (7.54%) 0.0067 (3.66%)
Full Model 182.54 1622.00 0.1833